Supposed a charge is at the origin

Determine the direction angle of the electric field at the point( 0.9, 1.2) measured counterclockwise from the positive x axis if the particle is positively charged?
1
Expert's answer
2021-09-09T10:50:56-0400

The electric field compotents

"E_x=k\frac{q}{(\sqrt{x^2+y^2})^3}x"

"E_y=k\frac{q}{(\sqrt{x^2+y^2})^3}y"

Hence,

"\tan\theta=\frac{E_y}{E_x}=\frac{y}{x}=\frac{1.2}{0.9}=1.33"

"\theta=\tan^{-1}1.33=53^{\circ}"